OpenStudy (redohawk):

Write an equation of a line perpendicular to y = 1 over 2x − 5 in slope-intercept form that passes through the point (2, −7). y= 1 over 2x − 7 y= 2x + 2 y= −2x − 5 y= −2x − 3

OpenStudy (redohawk):

@mathmale @MathLegend @mathmale @marissalovescats

OpenStudy (redohawk):

@hartnn @hartnn

OpenStudy (redohawk):

@hartnn @Hero

OpenStudy (redohawk):

@JungHyunRan

OpenStudy (redohawk):

@Abmon98 @beastieman21

OpenStudy (redohawk):

@goformit100

OpenStudy (redohawk):

anyone!!! @hartnn

hartnn (hartnn):

can you find the slope of y = 1 over 2x − 5 ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

slope of y =mx+c is just m

OpenStudy (anonymous):

slope of y =mx+c is just m

OpenStudy (redohawk):

you just copy and paste xD

OpenStudy (redohawk):

i have no clue, if i coul djust get the answer, that would be good ^^

OpenStudy (redohawk):

please! @hartnn

hartnn (hartnn):

remember that this is not an answering site, its a learning site , we don't give out answers here.

OpenStudy (redohawk):

ahhh... sorry :/

OpenStudy (redohawk):

i mean i know it is C or D right?

hartnn (hartnn):

yes now plug in this point in them (2, −7). x=2, and see for which option, y comes out to be -7

OpenStudy (redohawk):

its C coz -5 - 2 = -7

hartnn (hartnn):

its D y= −2x − 3 -2(2)-3 = -4-3 =-7